a train leaves at noon with constant velocity of 40 mi/hr another train leaves at 2pm with constant velocity of 70 mi/hr at what mile do they catch up?
Found 2 solutions by mananth, nerdybill:
Answer by mananth(16946)

Time difference = 2hours
catch up speed = 70-40=30 mph
distance at which they catch up = 30*2 = 60 miles

Answer by nerdybill(7384)

a train leaves at noon with constant velocity of 40 mi/hr another train leaves at 2pm with constant velocity of 70 mi/hr at what mile do they catch up?
.
Apply the distance formula of d=rt
.
Let x = time (hours) to catch up
then
40(x+2) = 70x
40x+80 = 70x
80 = 30x
80/30 = x
2.67 hours = x
or
2 hours and 40 minutes = x
.
2 pm + 2 hours and 40 minutes = 4:40 pm